OpenStudy (anonymous):

I'm having a little trouble with reducing fractions to the lowest term? I'll put the problem down!

OpenStudy (anonymous):

\[\frac{ 9 }{ 14 } \times -\frac{ 7 }{ 12 } \] now i know the 9 turns into a 3 and the 7 become 1 but the bottom half is confusing me. then the result is \[-\frac{ 3 }{ 8 }\] which is the answer. ._.

OpenStudy (anonymous):

\[\frac{ 3 }{ 14 }\times -\left(\begin{matrix}1 \\ 12\end{matrix}\right)\] Hopefully i'm writing this right!

OpenStudy (isaiah.feynman):

Not quite. Simply multiply 9 and 7 and multiply 14 and 12.

OpenStudy (anonymous):

9 x 7 = 63 14 x 12 = 168\[= -\frac{ 3 }{ 8 } \] ._. i should consult my math book haha.

OpenStudy (isaiah.feynman):

 See what I did??